An ideal gas at 10.3 °C and a pressure of 2.48 x 10^5 Pa occupies a volume of 3.12 m3. (a) How many moles of gas are present? (b) If the volume is raised to 5.45 m3 and the temperature raised to 25.6 °C, what will be the pressure of the gas?
1
Expert's answer
2020-09-25T09:09:47-0400

The ideal gas equation is given by

"pV=nRT" ….........(1)

where "p" is the pressure in "Pa"

"V" is the volume in "m^3"

n is the number of mole

"R = 8.314\ J.mol^{-1}.K^{-1}" is the universal gas constant and

"T" is the absolute temperature in "K" of the gas.


(a) Given, "p=2.48\times 10^5\ Pa,\ V=3.12\ m^3, T=10.3 \degree C=(273+10.3)K=283.3K"

From Eq.(1), "n=\frac{pV}{RT}"

Substituting the given values of "p,V,R,T", we get

"n=\frac{2.48\times 10^5\times 3.12}{8.314\times 283.3}\ mol"

"\therefore n=328.5\ mol"


(b) New volume "V'=5.45\ m^3"

New temperature "T'=25.6\degree C=(273+25.6)K=298.6K"

The new pressure is given by

"p'=\frac{nRT'}{V'}"

"p'=\frac{328.5\times 8.314\times 298.6}{5.45}\ Pa"

"\therefore p'=1.50\times 10^5\ Pa"


Answer: (a) There are 328.5 moles of gas.

(b) The new pressure of the gas will be "1.50\times 10^5\ Pa" .